# The fate of Schwarzschild–de Sitter black holes: nonequilibrium evaporation

### v1

@@ I Introduction
The discovery by Bekenstein and Hawking that black holes radiate thermally and carry entropy revolutionized our understanding of gravity and thermodynamics, revealing that classical event horizons are inherently quantum objects Bekenstein (1973, 1974); Bardeen et al. (1973); Hawking (1974, 1975). In an accelerating de Sitter universe, black holes coexist with a cosmological horizon, each characterized by its own surface gravity and associated temperature. Unlike asymptotically flat black holes, this dual-horizon structure places the system in intrinsic nonequilibrium: both horizons radiate, but generally at different temperatures. Despite significant progress, a complete understanding of how black holes evaporate in de Sitter space and how the second law of thermodynamics is realized in such spacetimes remains a fundamental open problem.
@@
The two-dimensional trace-anomaly method, pioneered by Polyakov Polyakov (1981) and developed in many subsequent works Christensen and Fulling (1977); Davies et al. (1976); Callan et al. (1992); Russo et al. (1992a); Fabbri and Navarro-Salas (2005), provides a powerful framework for semiclassical gravity. In this approach, the backreaction of $N$ conformal matter fields is encoded in an anomaly-induced effective action that captures quantum stress–energy effects in a geometrically transparent form, allowing controlled calculations of Hawking flux and backreaction in both asymptotically flat and cosmological settings (see, e.g., Callan et al. (1992); Russo et al. (1992b); Bousso and Hawking (1997); Balbinot and Fabbri (1999)). Yet an analytic, self-consistent treatment of Schwarzschild–de Sitter (SdS) evaporation that simultaneously tracks both horizons, real-time mass evolution, and the thermodynamic response of local observers has yet to be achieved.
@@
Recent studies have analyzed SdS and Kerr–de Sitter evaporation in four dimensions Bousso and Hawking (1998); Kanti and March-Russell (2003); Yoshida and Futamase (2004); Gregory et al. (2021), incorporating greybody factors and mode-by-mode fluxes for scalar, electromagnetic, and gravitational perturbations. These works demonstrate consistency with the generalized second law but rely on numerical computation and do not yield closed-form expressions for fluxes, mass evolution, or local thermodynamic quantities. A complementary, fully analytic framework can therefore provide valuable geometric insight and a benchmark for these numerical analyses.
@@
Here we construct such a platform using two-dimensional dilaton gravity with anomaly-induced backreaction. Starting from the spherical reduction of four-dimensional Einstein gravity with a cosmological constant, we obtain an exactly solvable two-dimensional model that reproduces the causal and thermodynamic structure of the SdS static patch. Incorporating the Polyakov effective action, we determine the anomaly-induced energy flux, derive the evolution equation for the black-hole mass, and compute local temperatures, energy densities, and fluxes measured by static observers. In two dimensions, the anomaly-induced action captures the complete semiclassical dynamics of the $s$-wave sector Alonso-Serrano et al. (2025), enabling a self-consistent treatment of backreaction.
@@
Building on this thermodynamic foundation, we develop a thermo-controlled estimate of the Page curve and show how quantum extremal surfaces and entanglement islands emerge naturally in the Unruh–de Sitter state. This connects the nonequilibrium thermodynamics of de Sitter black holes with their underlying information flow, providing a new analytic bridge between horizon thermodynamics and the island paradigm in a cosmological context.
These results yield a fully analytic, backreacted description of Schwarzschild–de Sitter evaporation that unifies semiclassical thermodynamics and information dynamics within a single, exactly solvable framework. The construction captures flux conservation, entropy balance, Page-like behavior, and endpoint classification under analytic control, complementing prior numerical analyses and offering exact geometric insight into evaporation, entropy production, and information recovery in multi-horizon spacetimes. Moreover, the methods developed here extend naturally to charged or regular (nonsingular) black holes Banks and O’Loughlin (1993); Trodden et al. (1993); Bogojevic and Stojkovic (2000); Easson and Brandenberger (2001); Easson (2003, 2018); Frolov and Zelnikov (2021); Davies et al. (2025), providing a foundation for future studies of quantum extremal surfaces and information flow in dynamical backgrounds.
@@ II From 4D Schwarzschild–de Sitter to 2D dilaton gravity
To begin, for completeness we briefly summarize the spherical reduction of four-dimensional Einstein–Hilbert (EH) gravity with cosmological constant to its two-dimensional dilaton form, following standard treatments and exhibit the Schwarzschild–de Sitter (SdS) solution in the reduced theory.  [^1] We closely follow the standard spherically reduced gravity (SRG) formalism Mann et al. (1990); Klosch and Strobl (1996); Grumiller et al. (2002).
@@
$U(X)=\frac{1}{2X},\qquad V(X)=\frac{1}{2}-\frac{\Lambda}{2}X\,.$ (6)
@@ II.1 Weyl frame and SdS solution
$\tilde{V}(X):=e^{Q(X)}V(X)=X^{-1/2}\!\left(\tfrac{1}{2}-\tfrac{\Lambda}{2}X\right).$ (9)
@@
$w^{\prime}(X)=\tilde{V}(X),\qquad\xi(X)=e^{Q(X)}\big[w(X)-2G_{4}M\big],$ (10)
@@
Integration of Eq. 10 gives
$\begin{split}w(X)&=\!\int^{X}\!\tilde{V}(y)\,dy=\!\int\!\Big(\tfrac{1}{2r}-\tfrac{\Lambda}{2}r\Big)2r\,dr\\ &=\,r-\tfrac{\Lambda}{3}r^{3}+{\rm const}.\end{split}$ (12)
@@ III.1 Conserved flux and mass–loss law
Since $\chi=\partial_{t}=\partial_{v}$ in the stationary patch, the stress–tensor components coincide,
$T^{r}{}_{t}=T^{r}{}_{v}=\frac{\mathcal{J}}{IX^{\prime}}=\frac{\mathcal{J}}{2r^{2}},$ (19)
so $\mathcal{J}$ represents the single conserved Killing energy flux across the static patch.
@@
directed outward for $\kappa_{b}>\kappa_{c}$. Identifying $M$ with the total energy inside the static patch, the evolution follows
@@ IV.1 Flux, equilibrium and evolution timescales
only the first regime occurs in the neutral black hole case; charged (Reissner-Nordström) or modified-gravity extensions may realize the others. We now analyze these equilibria in detail and identifies the Nariai configuration as the unique zero-flux limit.
@@ V Quasi-stationary geometry from localized Polyakov action
To describe this quasi-stationary configuration from first principles, we solve the field equations derived from the localized Polyakov action. The resulting geometry represents a fully backreacted, adiabatic steady state of the Unruh–de Sitter type, in which the flux $\mathcal{J}$ is constant across the patch and mediates a net flow of energy from the black hole to the cosmological horizon. This solution demonstrates that the nonequilibrium evaporation dynamics of SdS in the Unruh–de Sitter state emerge self-consistently from the Einstein equations in two-dimensional dilaton gravity. In this analysis, all leading quantum effects are incorporated through the large-$N$, one-loop treatment of $N$ conformal fields.
@@ V.2 Solution structure and horizon regularity
This same anomaly–induced backreaction that governs the semiclassical dynamics of the SdS patch also plays a decisive role in other multi-horizon systems. An analogous mechanism destabilizes the inner (Cauchy) horizons of charged and rotating black holes, where the trace anomaly is the minimal quantum enforcer of strong cosmic censorship Easson (2025) by converting inner horizons into null curvature singularities Hollands et al. (2020).
@@ VI.1 The Nariai limit and zero–flux equilibrium
In our two–dimensional dilaton gravity description, the Nariai state is realized when the effective potential satisfies
@@ VIII.1 Entropy balance and first-law structure
$\delta M=T_{b}\,\delta S_{b}-T_{c}\,\delta S_{c},$ (45)
where the minus sign reflects the opposite Killing orientation of the cosmological horizon. This “first law” describes variations among nearby static configurations and should not be interpreted as a dynamical evolution equation. The time-dependent semiclassical evolution is governed instead by the conserved energy flux.
@@ A.5 Dimensional-reduction anomaly
The present analysis employs the standard spherically reduced two–dimensional dilaton–gravity action augmented by the Polyakov term, which correctly reproduces the trace anomaly of $N$ conformal matter fields but omits the so–called dimensional‐reduction anomaly: additional state‐dependent terms that arise when integrating out angular modes in the four–dimensional theory. These corrections can modify the quantitative relation between the 4D and 2D fluxes, rescaling the magnitude of $\mathcal{J}$ or shifting the effective dilaton potential. However, they do not affect the sign structure of the flux or the inequalities central to our conclusions. The positivity of $\Delta=\kappa_{b}^{2}-\kappa_{c}^{2}$, the direction of energy flow $\mathcal{J}\propto\Delta(M)$, and the non–negativity of the generalized entropy production rate $\dot{S}_{\mathrm{gen}}\sim\mathcal{J}(1/T_{c}-1/T_{b})$ follow purely from geometric relations among the Schwarzschild–de Sitter horizons and thus remain unchanged under any overall normalization of the stress tensor. In this sense, the dimensional–reduction anomaly alters how fast evaporation proceeds but not which way it goes or whether the generalized second law holds. This line of thinking is further justified by results from the four dimensional case Hollands et al. (2020).

### v2

@@ I Introduction
The discovery by Bekenstein and Hawking that black holes radiate thermally and carry entropy revolutionized our understanding of gravity and thermodynamics, revealing that classical event horizons are inherently quantum objects Bekenstein (1973, 1974); Bardeen et al. (1973); Hawking (1974, 1975). In an accelerating de Sitter universe, black holes coexist with a cosmological horizon, each characterized by its own surface gravity and associated temperature. Unlike asymptotically flat black holes, this dual-horizon structure places the system in intrinsic nonequilibrium: both horizons radiate, but generally at different temperatures. Despite significant progress, a complete understanding of black-hole evaporation in de Sitter space and of how the second law of thermodynamics is realized in such spacetimes remains a fundamental open problem.
@@
A powerful framework for addressing semiclassical backreaction is provided by the two-dimensional trace-anomaly method pioneered by Polyakov Polyakov (1981) and developed in many subsequent works Christensen and Fulling (1977); Davies et al. (1976); Callan et al. (1992); Russo et al. (1992a); Fabbri and Navarro-Salas (2005). In this approach, the backreaction of $N$ conformal matter fields is encoded in an anomaly-induced effective action that captures quantum stress–energy effects in a geometrically transparent form, allowing controlled calculations of Hawking flux and backreaction in both asymptotically flat and cosmological settings (see, e.g., Callan et al. (1992); Russo et al. (1992b); Bousso and Hawking (1997); Balbinot and Fabbri (1999)). Yet an analytic, self-consistent treatment of Schwarzschild–de Sitter (SdS) evaporation that simultaneously tracks both horizons, real-time mass evolution, and the thermodynamic response of local observers has not yet been achieved.
Two-dimensional dilaton gravity has long served as a laboratory for exploring black-hole thermodynamics, semiclassical backreaction, and departures from classical singular behavior, including early studies of Hawking radiation, nonsingular geometries, and regular Schwarzschild and Schwarzschild–de Sitter solutions Poisson and Israel (1990); Banks and O’Loughlin (1993); Trodden et al. (1993); Bogojevic and Stojkovic (2000); Easson and Brandenberger (2001); Easson (2003, 2018); Frolov and Zelnikov (2021); Davies et al. (2025). While these works provide important conceptual and technical insights into regularity, backreaction, and effective descriptions of black-hole interiors, they do not incorporate the full SdS static patch or yield a closed-form, globally backreacted, two-horizon evaporation solution.
@@
In parallel, four-dimensional studies of Schwarzschild–de Sitter and Kerr–de Sitter evaporation have analyzed Hawking emission using greybody factors and mode-by-mode fluxes for scalar, electromagnetic, and gravitational perturbations Bousso and Hawking (1998); Kanti and March-Russell (2003); Yoshida and Futamase (2004); Gregory et al. (2021), demonstrating consistency with the generalized second law. Related semiclassical analyses have constructed perturbatively self-consistent SdS solutions by solving backreaction equations for quantum fields on a fixed background Akhmedov et al. (2024). However, these approaches rely on numerical computation or perturbative control and do not yield closed-form expressions for the evaporation fluxes, mass evolution, or local thermodynamic observables.
@@
Here we construct an analytic framework that addresses these limitations. Starting from the spherical reduction of four-dimensional Einstein gravity with a cosmological constant, we obtain an exactly solvable two-dimensional dilaton gravity model that reproduces the causal and thermodynamic structure of the SdS static patch. Incorporating the Polyakov effective action, we determine the anomaly-induced energy flux, derive the evolution equation for the black-hole mass, and compute local temperatures, energy densities, and fluxes measured by static observers. In two dimensions, the anomaly-induced action captures the complete semiclassical dynamics of the $s$-wave sector Polyakov (1981); Russo et al. (1992a); Alonso-Serrano et al. (2025), enabling a self-consistent treatment of backreaction.
@@
Quantum extremal surfaces (QES) and entanglement islands in de Sitter–like settings have been analyzed previously in simplified two-dimensional models, most prominently in the near-Nariai JT limit Aalsma et al. (2019); Aalsma and Sybesma (2021) and in related treatments of cosmological or near-extremal throats Hartman et al. (2020); Shaghoulian and Susskind (2022), as well as in studies of vacuum-dependent evaporation and island formation in near-extremal RNdS or JT-like reductions Bhattacharjee et al. (2021). These works exploit the analytic tractability of near-horizon geometries, where the dilaton takes a simple form and the generalized entropy can be extremized explicitly.
Our framework differs in several essential respects. We employ the full spherical reduction of four-dimensional SdS and incorporate anomaly-induced backreaction throughout the entire static patch, rather than restricting to a near-horizon JT limit. In our Eddington–Finkelstein formulation, the Unruh–de Sitter (UdS) state is implemented through constant state functions $t_{\pm}$ (equivalently the conserved flux $\mathcal{J}$), so the backreacted solution satisfies the UdS regularity conditions without requiring an additional state-defining coordinate transformation. Consequently, our discussion of islands emphasizes the thermodynamic necessity of the Page transition in a global nonequilibrium steady state, rather than a microscopic extremization of $S_{\rm gen}$ in a near-horizon throat. A detailed QES extremization in the fully backreacted SdS geometry is left for future work.
@@ II From 4D Schwarzschild–de Sitter to 2D dilaton gravity
To fix notation and for completeness, we briefly summarize the spherical reduction of four-dimensional Einstein–Hilbert gravity with a cosmological constant to its two-dimensional dilaton formulation, following standard treatments, and recover the Schwarzschild–de Sitter (SdS) solution within the reduced theory. [^1] We closely follow the standard spherically reduced gravity (SRG) formalism Mann et al. (1990); Klosch and Strobl (1996); Grumiller et al. (2002).
@@
$U(X)=\frac{1}{2X},\qquad V(X)=1-\Lambda X.$ (6)
@@ II.1 Weyl frame and SdS solution
$\tilde{V}(X):=e^{Q(X)}V(X)=X^{-1/2}\!\left(1-\Lambda X\right).$ (9)
@@
$w^{\prime}(X)=\frac{1}{2}\,\tilde{V}(X),\qquad\xi(X)=e^{Q(X)}\big[w(X)-2G_{4}M\big],$ (10)
@@
Integration of (10) gives
$\begin{split}w(X)&=\int^{X}\frac{1}{2}\tilde{V}(y)\,dy=\int\frac{1}{2}\Big(\tfrac{1}{r}-\Lambda r\Big)2r\,dr\\ &=r-\tfrac{\Lambda}{3}r^{3}+{\rm const}.\end{split}$ (12)
@@ III Semiclassical trace–anomaly backreaction
Note that in contrast to near-Nariai JT gravity analyses, where an explicit closed-form solution for the dilaton evaluated in the Unruh–de Sitter state is required in order to compute the generalized entropy exactly, our analysis does not rely on an explicit UdS–state dilaton profile. Instead, we work directly with the spherically reduced dilaton gravity model describing the full Schwarzschild–de Sitter static patch and determine the semiclassical geometry and stress tensor in the UdS state via the anomaly-induced effective action and Killing energy conservation. This is sufficient to fix the stationary Hawking flux, local thermodynamic observables, and entropy production without requiring an explicit analytic expression for the dilaton as a function of spacetime coordinates in the UdS state.
@@ III.1 Conserved flux and mass–loss law
The conserved Killing energy current is $j^{\mu}=-T^{\mu}{}_{\nu}\chi^{\nu}$ with $\chi=\partial_{t}$. In ingoing EF coordinates one has $\chi=\partial_{v}$, so the radial Killing flux is $j^{r}=-T^{r}{}_{v}$, yielding
$T^{r}{}_{v}=\frac{\mathcal{J}}{IX^{\prime}}=\frac{\mathcal{J}}{2r^{2}}.$ (19)
@@
Here $M$ denotes the two–dimensional Casimir mass parameter defined by $\mathcal{C}=2G_{4}M$ (up to an additive constant), while the corresponding four–dimensional physical mass is $M_{\rm phys}=M/(2G_{4})$.
@@
directed outward for $\kappa_{b}>\kappa_{c}$. Identifying $M$ with the Casimir mass parameter controlling the SdS geometry, the evolution follows
@@ III.2 Local temperatures and Tolman redshift
The quantities $T_{h}=\kappa_{h}/2\pi$ appearing in the First Law (45) are the standard Killing temperatures, defined by the surface gravities of the two horizons. Local static observers, however, measure Tolman-redshifted temperatures $T_{\rm loc}(r)=T_{h}/\sqrt{f(r)}$, which differ from $T_{h}$ except in special equilibrium cases (e.g. the Nariai limit). Because de Sitter space lacks a preferred asymptotic region, the distinction between Killing, Tolman, and other proposed “global” temperature notions (e.g. the one advocated in Volovik (2023)) is conceptually important.
In the present work, this ambiguity does not affect the evaporation dynamics: the mass-loss law is controlled entirely by the conserved Killing energy flux $\mathcal{J}$, which is a radially invariant quantity and therefore independent of which temperature convention one adopts. The horizon Clausius relations and the generalized second law follow directly from $\mathcal{J}$ and the geometric relation between $\kappa_{b}$ and $\kappa_{c}$, rather than from any particular observer-adapted temperature definition. For recent discussions of observer-dependent thermodynamics in multi-horizon Reissner–Nordström–de Sitter spacetimes, see Aalsma et al. (2025).
@@ IV.1 Flux, equilibrium and evolution timescales
only the first regime occurs in the neutral black hole case. Charged (Reissner–Nordström–de Sitter) black holes or modified-gravity extensions are known to admit interior solutions with $\kappa_{b}=\kappa_{c}\neq 0$ (“lukewarm” configurations), and may therefore realize the latter two regimes, but a detailed analysis of such cases lies beyond the scope of this paper.
We now analyze these equilibria in detail and identifies the Nariai configuration as the unique zero-flux limit.
@@ V Quasi-stationary geometry from localized Polyakov action
To describe this quasi-stationary configuration from first principles, we solve the field equations derived from the localized Polyakov action. The resulting geometry represents a fully backreacted, adiabatic steady state of the Unruh–de Sitter type, in which the flux $\mathcal{J}$ is constant across the static patch and mediates a net flow of energy from the black hole to the cosmological horizon. This solution demonstrates that the nonequilibrium evaporation dynamics of Schwarzschild–de Sitter spacetime in the Unruh–de Sitter state emerge self-consistently from the (dimensionally reduced) Einstein equations. In this analysis, all leading quantum effects are incorporated through a large-$N$, one-loop treatment of $N$ conformal fields.
@@ V.2 Solution structure and horizon regularity
This same anomaly–induced backreaction that governs the semiclassical dynamics of the SdS patch also plays a decisive role in other multi-horizon systems. An analogous mechanism destabilizes the inner (Cauchy) horizons of charged and rotating black holes, where the trace anomaly is the minimal quantum enforcer of strong cosmic censorship by converting inner horizons into null curvature singularities Birrell and Davies (1978); Poisson and Israel (1990); Ori (1991); Brady and Smith (1995); Hollands et al. (2020a, b); Cardoso et al. (2018); Dafermos (2005); Easson (2025).
@@ VI.1 The Nariai limit and zero–flux equilibrium
In our dilaton gravity description, the Nariai state is realized when the effective potential satisfies
@@ VIII.1 Entropy balance and first-law structure
$2\delta M=T_{b}\,\delta S_{b}-T_{c}\,\delta S_{c},$ (45)
where the minus sign reflects the opposite Killing orientation of the cosmological horizon. [^6] Different conventions exist for defining temperatures and entropies in 2D dilaton gravity. Some authors define the Wald entropy as $S=2\pi X$, which leads to a first law of the form $\delta M=2T\,\delta S$, whereas in the spherical-reduction normalization used here $S=X/(4G_{2})$ (equivalently $S=\pi r^{2}/G_{4}$) reproduces the standard 4D area law and yields $\delta M=T_{b}\,\delta S_{b}-T_{c}\,\delta S_{c}$ with no extra factor. Similarly, the temperatures $T_{h}=\kappa_{h}/2\pi$ are the Killing temperatures that enter the geometric first law, while local static observers measure Tolman-redshifted values $T_{\rm loc}=T_{h}/\sqrt{f(r)}$. This distinction does not affect the mass-loss law, which depends only on the conserved Killing flux $\mathcal{J}$. This “first law” describes variations among nearby static configurations and should not be interpreted as a dynamical evolution equation. The time-dependent semiclassical evolution is governed instead by the conserved energy flux.
@@ IX.2.2 Cosmological counterpart
While cosmological- and black-hole–side quantum extremal surfaces have been identified in equilibrium and near-horizon models, the role of a conserved, anomaly-induced nonequilibrium flux in selecting the physically relevant island saddle for a given observer in the full SdS static patch has not been explicitly analyzed previously.
@@ A.5 Dimensional-reduction anomaly
The present analysis employs the standard spherically reduced two–dimensional dilaton–gravity action augmented by the Polyakov term, which correctly captures the structure of the trace anomaly of $N$ conformal matter fields but omits the so–called dimensional‐reduction anomaly: additional state‐dependent terms that arise when integrating out angular modes in the four–dimensional theory. These corrections can modify the quantitative relation between the 4D and 2D fluxes, rescaling the magnitude of $\mathcal{J}$ or shifting the effective dilaton potential. However, they do not affect the sign structure of the flux or the inequalities central to our conclusions. The positivity of $\Delta=\kappa_{b}^{2}-\kappa_{c}^{2}$, the direction of energy flow $\mathcal{J}\propto\Delta(M)$, and the non–negativity of the generalized entropy production rate $\dot{S}_{\mathrm{gen}}\sim\mathcal{J}(1/T_{c}-1/T_{b})$ follow purely from geometric relations among the Schwarzschild–de Sitter horizons and thus remain unchanged under any overall normalization of the stress tensor. In this sense, the dimensional–reduction anomaly alters how fast evaporation proceeds but not which way it goes or whether the generalized second law holds. This line of thinking is further justified by results from the four dimensional case Hollands et al. (2020a).